# Tiny Time Mixers (TTMs): Fast Pretrained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series

### 11 Jan 2024 v2

@@ 4.3 Computational Benefits of TTM
Table  2 compares the computational benefits of TTM with GPT4TS, the most popular LLM based TS model and the current best SOTA in few-shot setting. We run both models in a single A100 GPU with matching batch sizes for a fair comparison. Multi-GPU usage is deliberately avoided in this experiment to eliminate IPC overheads and ensure precise metric calculations. Since GPT4TS is based on GPT-2, it consumes a huge model footprint and exhibits slow execution as compared to TTM. In specific, TTM achieves 14X cut in Finetune (FT) parameters (NPARAMS) and a remarkable 106X reduction in Total (TL) parameters. This reduction in model footprint further leads to a significant reduction in fine-tuning EPOCH TIME by 65X, MAX MEMORY usage by 27X, and total inference time on the entire test data (TEST TIME) by 54X. The computational advantages also apply over other LLM-based time series models like LLMTime, built on LLaMA, which is larger than GPT-2 (utilized in GPT4TS). In fact, LLMTime used a very small test set on these datasets as opposed to the standard test set to overcome this slow-execution behavior.
4.4 TTM’s effectiveness in cross-channel and exogenous/control modeling
Since the datasets ($D1$) used in previous experiments do not have exogenous variables, we evaluate the effectiveness of TTM on 4 other datasets ($D2$, as explained in  4.1) to quantify its benefits. Table 5 shows the performance of the Monash-pretrained TTM model fine-tuned on the target data with exogenous mixer module and decoder channel-mixing enabled (TTM-CM). We compare the full-shot performance of TTM-CM with the pretrained TTM used in zero-shot setting (TTM-Zero-shot), and plain TTM (TTM) and other primary SOTAs (PatchTST, TSMixer variants, and GPT4TS) trained from scratch. Remaining SOTAs are not reported in Table 5 considering their inferior performance and space constraints. Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with Cross-Channel reconciliation head (TSMixer-CC) Ekambaram et al. (2023) as they are latest SOTAs in channel-correlation modeling. We can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating that the power of TTM in capturing inter-channel correlations that are prevalent in purely multivariate datasets.
@@ 4.5 Ablation Studies
Effect of Pretraining: Table 6 illustrates the advantages of the proposed pretraining (PT) approach in comparison to randomly initialized (RI) model weights. In the case of zero-shot, the pretrained TTM (PT) exhibits 36% improvement over RI. This outcome is expected, as random weights are directly employed for forecasting with RI. For the 5% few-shot scenario, PT achieves noteworthy improvements of 21% and 12% over RI when fine-tuned for 5 and 50 epochs, respectively. This underscores the utility of pretrained weights in facilitating quick learning with a limited amount of target data. Even in the case of 10% few-shot, PT continues to demonstrate an improved performance of 9% for 5 epochs and 2% for 50 epochs. Thus, the performance impact of leveraging pretrained weights greatly increases as the size of the training data and the available time-to-finetune reduces.
Effect of Adaptive Patching: The Adaptive patching presented in Section 3.1.1 is a key component of the TTM backbone. Table 7 demonstrates the relative improvements in MSE when we employ adaptive patching with respect to a vanilla TTM backbone without this module. We can see that adaptive patching adds an extra 4% improvement during zero-shot and 2% improvement for 10% few-shot. MSE improvements are observed in almost all datasets which justifies the effectiveness of the adaptive patching module when evaluated on multiple datasets of different resolutions.
Effect of Augmentation via Downsampling: Table 8 compares the zero-shot test performances between the TTM model trained on the original Monash data and the TTM trained on the augmented Monash data after downsampling as described in Section 3.1. It is evident that there is a significant improvement of 30% for the model pretrained on the augmented datasets as it adds more data coverage across various resolutions. This observation highlights the strength of the proposed TTM model to learn from diverse datasets together, and increasing the number of datasets for each resolution helps.
@@ 5 Conclusions and Future Work
Considering the diversity and limited availability of public TS pretraining data, pretraining large models for effective transfer learning poses several challenges in time series. Hence, we propose TTM, a multi-level Tiny Time Mixer model designed for efficient pretraining on limited diverse multi-resolution datasets. TTM achieves state-of-the-art results in zero/few-shot forecasting, offering significant computational efficiency and supporting cross-channel and exogenous variables — critical features lacking in existing popular methods. Going forward, we plan to extend our approach to many other downstream tasks beyond forecasting for a purely foundational approach in time series.

### 17 Jan 2024 v3

@@ 4.3 Computational Benefits of TTM
Table  2 compares the computational benefits of TTM over GPT4TS, the most popular LLM-TS model and the current best SOTA in few-shot setting. For a fair comparison, we run both models with the best-reported parameters in a single A100 GPU with the same setup (Details in Appendix). Since GPT4TS is based on GPT-2, it consumes a huge model footprint and exhibits slow execution. In specific, TTM achieves 14X cut in Fine-tune (FT) parameters (NPARAMS) and a remarkable 106X reduction in Total (TL) parameters. This reduction in model footprint further leads to a significant reduction in fine-tuning EPOCH TIME by 65X, MAX MEMORY usage by 27X, and total inference time on the entire test data (TEST TIME) by 54X. The computational benefits also apply over other LLM-based time series models like LLMTime, built on LLaMA, which is larger than GPT-2. In fact, LLMTime used a very small test set on these datasets as opposed to the standard test set to overcome this slow-execution.
4.4 TTM’s effectiveness in cross-channel and exogenous modeling
Since the datasets ($D1$) used in previous experiments do not have exogenous variables, we evaluate the effectiveness of TTM on 4 other datasets ($D2$, as explained in  4.1) to quantify its benefits. Since these datasets are already very small, we used their full data for fine-tuning. Table 5 shows the performance of the pre-trained TTM model fine-tuned on the target data with exogenous mixer module and decoder channel-mixing enabled (TTM-CM). We compare the performance of TTM-CM with the pre-trained TTM used in zero-shot setting (TTM-Zero-shot), and plain TTM (TTM) and other primary SOTAs (PatchTST, TSMixer variants, and GPT4TS) trained from scratch. Other SOTAs are not reported here, considering their inferior performance and space constraints. Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with Cross-Channel reconciliation head (TSMixer-CC) Ekambaram et al. (2023) as they are the latest SOTAs in channel-correlation modeling. We can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating the power of TTM in capturing inter-channel correlations.
@@ 4.5 Ablation Studies
Effect of Pre-training: Table 6 illustrates the advantages of the proposed pre-training (PT) approach in comparison to randomly initialized (RI) model weights. In the case of zero-shot, the pre-trained TTM (PT) exhibits 36% improvement over RI. This outcome is expected, as random weights are directly employed for forecasting with RI. For the 5% few-shot scenario, PT achieves noteworthy improvements of 21% and 12% over RI when fine-tuned for 5 and 50 epochs, respectively. This underscores the utility of pre-trained weights in facilitating quick learning with a limited amount of target data. Even in the case of 10% few-shot, PT continues to demonstrate an improved performance of 9% for 5 epochs and 2% for 50 epochs. Thus, the performance impact of leveraging pre-trained weights greatly increases as the size of the training data and the available time-to-fine-tune reduces.
Effect of Adaptive Patching: The Adaptive patching presented in Section 3.1.1 is a key component of the TTM backbone. Table 7 demonstrates the relative improvements when we employ adaptive patching with respect to a vanilla TTM backbone without this module. We can see that adaptive patching adds an extra 4% improvement during zero-shot and 2% improvement for 10% few-shot. Performance improvements are observed in almost all datasets which justifies the effectiveness of the adaptive patching module when evaluated on multiple datasets of different resolutions.
Effect of Augmentation via Downsampling: Table 8 compares the zero-shot test performances between the TTM model trained on the original Monash data and the TTM trained on the augmented Monash data after downsampling as described in Section 3.1. There is a significant improvement of 30% for the model pre-trained on the augmented datasets as it adds more data coverage across various resolutions. This observation highlights the strength of the proposed TTM model to learn from diverse datasets together, and increasing the number of datasets for each resolution helps.
@@ 5 Conclusions and Future Work
Considering the diversity and limited availability of public TS pre-training data, pre-training large models for effective transfer learning poses several challenges in time series. Hence, we propose TTM, a multi-level Tiny Time Mixer model designed for efficient pre-training on limited diverse multi-resolution datasets. TTM achieves state-of-the-art results in zero/few-shot forecasting, offering significant computational efficiency and supporting cross-channel and exogenous variables — critical features lacking in existing popular methods. Going forward, we plan to extend our approach to many other downstream tasks beyond forecasting for a purely foundational approach in time series.